OpenStudy (anonymous):

Another difficult question!! (x+6)/10 divided by (36-x^2)/5

myininaya (myininaya):

a/b divided by c/d= a/b times d/c

OpenStudy (anonymous):

I did set up the reciprocal so I am multiplying but that is where I get stuck!

myininaya (myininaya):

so you have [(x+6)*5]/[10*(36-x^2)] correct?

OpenStudy (anonymous):

Yes

myininaya (myininaya):

10=2*5 so a 5 can cancel?

OpenStudy (anonymous):

so (36-x^2) is equal to (6-x)(6+x)

OpenStudy (anonymous):

Ok

myininaya (myininaya):

since you have 6+x on top and 6+x on bottom that cancels

OpenStudy (anonymous):

right

myininaya (myininaya):

so you 1/[2*(6-x)]

OpenStudy (anonymous):

I cancelled the 6+x and now have 5/10(6-x)

myininaya (myininaya):

but what did we say about that 10

OpenStudy (anonymous):

Oh yeah, missed that!!

myininaya (myininaya):

ok cool!

OpenStudy (anonymous):

so then shall I multiply the 2(6-x) and get 12-2x

myininaya (myininaya):

if you want. When you take calculus though it will be better to leave it as the first way, but if your teacher wants you to mutliply the bottom thats correct

OpenStudy (anonymous):

Ok, I am not really sure how he wants it, I am taking an online math class and it sucks!!

myininaya (myininaya):

the test multiple choice?

OpenStudy (anonymous):

No, you just type your answer in

myininaya (myininaya):

ok i think you can type your answer either way. it should take it either way unless you can't use grouping symbols in your solution

OpenStudy (anonymous):

I will try. Thank you for your help!!

myininaya (myininaya):

np